# General Order Virtual Element Methods for Neumann Boundary Optimal Control Problems in Saddle Point Formulation

### v1

@@ Abstract.
In this work, we explore the application of the Virtual Element Methods for Neumann boundary Optimal Control Problems in saddle point formulation. The method is proposed for arbitrarily polynomial order of accuracy and general polygonal meshes. Our contribution includes a rigorous a priori error estimate that holds for general polynomial degree. On the numerical side, we present (i) an initial convergence test that reflects our theoretical findings, and (ii) a second test case based on a more application-oriented experiment. For the latter test, we focus on the role of VEM stabilization, conducting a detailed experimental analysis, and proposing an alternative structure-preserving strategy to circumvent issues related to the choice of the stabilization parameter.
1. Introduction
@@
In this paper, we consider Virtual Element Methods (VEM) approximation for optimal Neumann boundary control. Optimal Control Problems (OCPs) constrained to partial differential equations (PDEs) are widespread in many practical applications in scientific and engineering contexts, see e.g. [2, 30, 32, 33] and the references therein. The main goal of OCPs is to reach a beneficial behavior of the PDE solution by employing an external variable, the control. The role of the control consists in making the state solution similar to a desired configuration using constrained minimization of a functional. The control is distributed when its action is defined all over the domain, while we talk about boundary control if the control affects the boundary conditions. Due to its coupled nature, the numerical analysis for OCPs constrained to PDEs is a fascinating mathematical challenge. Several works have been proposed with Finite Element Methods, Discontinuous Galerkin Methods, Finite Volumes and hybrid methods: we refer the reader to this far from exhaustive list and the reference therein [18, 21, 24, 27, 42, 43, 44]
@@
Lately, great attention has been paid to VEM, a family of polygonal methods for simulating PDEs. VEM can be applied on arbitrarily polygonal meshes, guaranteeing huge flexibility in terms of geometrical complexity and they might play a crucial role in many real-life scenarios. Indeed, after being conceived in the seminal paper [3], they have been applied to various contexts of interest [5, 7, 13, 19, 20, 23, 41]. The investigation of boundary OCPs through VEM is still limited yet of utmost importance. Indeed, the versatility of arbitrary polygonal meshes, combined with the action of control strategies on complex geometries, might be not only a challenging theoretical topic but also relevant from a practical viewpoint. To our knowledge, $C^{1}$-VEM is presented in [16] for distributed OCPs constrained to elliptic equations and with a constrained state variable. The VEM discretization for distributed control is the topic of several investigations, see e.g. [34, 36, 37, 38, 40]. Some VEM works related to boundary OCPs are [39] for the Dirichlet case with control constraints and [25] for the Neumann case with control constraints and approximation of the first order.
In this contribution, we focus on the unconstrained Neumann boundary control problem in a saddle point structure. This naturally derives from the optimality conditions and considers the coupled nature of the system leading to more robust and stable systems solving for all the variables at once. This formulation is generally more effective than iterative algorithms. To the best of our knowledge, we differentiate from previous contributions in the following terms:
  • the presented method has arbitrary polynomial accuracy;
@@
  • we analyze the problem in saddle point formulation as presented in [28];
  • we prove well-posedness and a priori error estimates for a general polynomial degree and variable coefficients;
  • numerical tests cover high-order schemes;
  • we numerically investigate how the stabilization parameter affects the method’s accuracy.
Moreover, we propose a structure-preserving strategy to circumvent issues related to VEM stabilizations, based on recent developments in VEM theory [8, 9, 10, 12].
@@
The paper is structured as follows: in Section 2 we introduce the linear-quadratic Neumann boundary OCP at the continuous level in a saddle point framework. Section 3 focuses on the VEM approximation of such optimization problem and proves the well-posedness of the saddle point structure. In Section 4 we derive optimal a priori error estimates. Section 5 presents two test cases: a convergence analysis test and a more application-oriented test for which we discuss the role of the stabilization parameter, proposing a workaround strategy using structure-preserving VEM. Finally, conclusions follow in Section 6.
@@
2. Problem Formulation
Let us consider $\Omega\subset\mathbb{R}^{2}$ as an open and bounded regular domain. We denote as $\Gamma_{D}$ and $\Gamma_{N}$ the portions of the boundary $\partial\Omega$ where Dirichlet and Neumann conditions apply. Moreover, we define a control boundary $\Gamma_{C}$, where the control acts in form of Neumann condition. We remark that $\Gamma_{D}\cup\Gamma_{N}\cup\Gamma_{C}=\partial\Omega$ and $\Gamma_{D}\cap\Gamma_{N}\cap\Gamma_{C}=\emptyset$. Moreover, for simplicity we assume $\Gamma_{D}\neq\emptyset$.
@@ 2. Problem Formulation
The case of non-homogeneous boundary conditions can be tackled by standard lifting procedures [31]. Given a desired state $y_{\mathrm{d}}\in Y_{\text{d}}=L^{2}(\Omega_{\text{obs}})$, with $\Omega_{\text{obs}}\subset\Omega$, a forcing term $f\in L^{2}(\Omega)$ and a penalization parameter $\alpha>0$ on the control action, we solve the following problem.
@@
(1) $\begin{cases}-\nabla\cdot(\mathcal{K}\nabla w)+\gamma w=f&\text{in }{\Omega},\\ \displaystyle w=0&\text{on }{\Gamma_{D}},\\ \displaystyle\frac{\partial w}{\partial n}=v&\text{on }{\Gamma_{C}},\vspace{0.5mm}\\ \displaystyle\frac{\partial w}{\partial n}=0&\text{on }{\Gamma_{N}},\\ \end{cases}$
where $\mathcal{K}$ and $\gamma$ are a uniformly elliptic tensor valued function and a scalar function, respectively, both considered sufficiently smooth.
@@
(2) ${a(y,q)}=\left(u,q\right)_{\Gamma_{C}}+\left(f,q\right)_{\Omega}\quad\forall q\in Y,$
@@
(3) $a(y,q)\vcentcolon=\left(\mathcal{K}\nabla y,\nabla q\right)_{\Omega}+\left(\gamma y,q\right)_{\Omega}\,\quad\forall y,q\in Y.$
Since $a(\cdot,\cdot)$ is continuous and coercive and $\left(u,q\right)_{\Gamma_{C}}$ is continuous, (2) is well-posed [31]. Let $\left\lvert\left\lvert\left\lvert\cdot\right\rvert\right\rvert\right\rvert_{Y}$ denote the norm induced by ${a}$, i.e. for each $w\in Y$
(4) $\left\lvert\left\lvert\left\lvert w\right\rvert\right\rvert\right\rvert_{Y}^{2}{\;\vcentcolon=\;}\left\lVert{\sqrt{\mathcal{K}}\nabla w}\right\rVert_{\mathrm{L}^{2}\!(\Omega)}^{2}+\left\lVert{\sqrt{\gamma}w}\right\rVert_{\mathrm{L}^{2}\!(\Omega)}^{2}\,.$
@@
For any $w\in Y$, there exists $C>0$ independent of $h$ such that
(5) $\left\lVert{w}\right\rVert_{\mathrm{H}^{1}\!(Y)}\leq C\left\lvert\left\lvert\left\lvert w\right\rvert\right\rvert\right\rvert_{Y}$
holds true, due to a Poincarè inequality.
@@
(7) $\mathscr{L}(y,u,p)=J(y,u)+\displaystyle a(y,p)-\left(u,p\right)_{\Gamma_{C}}-\left(f,p\right)_{\Omega}.$
@@
(8) $\begin{cases}D_{y}\mathscr{L}(y,u,p)[z]=0&\forall z\in Y,\\ D_{u}\mathscr{L}(y,u,p)[v]=0&\forall v\in U,\\ D_{p}\mathscr{L}(y,u,p)[q]=0&\forall q\in Y,\\ \end{cases}$
where $D_{y},D_{u}$ and $D_{p}$ denotes the differentiation with respect to the three variables, respectively. In this way, the constrained optimization problem is reframed in an unconstrained optimization and the minimum of (6) is represented by the state and control components of the stationary point of (7). The strong formulation of the optimality system reads:
@@
(9) $\begin{cases}\displaystyle y\boldsymbol{\mathbbm{1}}_{\{\Omega_{\text{obs}}\}}-\nabla\cdot(\mathcal{K}\nabla p)+\gamma p=y_{\mathrm{d}}\boldsymbol{\mathbbm{1}}_{\{\Omega_{\text{obs}}\}}&\text{ in }\Omega,\quad\text{{adjoint equation}}\\ \alpha u-p=0&\text{ in }\Gamma_{C},\;\;\text{{optimality condition}}\\ -\nabla\cdot(\mathcal{K}\nabla y)+\gamma y=f&\text{ in }\Omega,\quad\text{{state constraint}}\\ \displaystyle\frac{\partial y}{\partial n}=u\quad\text{and}\quad\frac{\partial p}{\partial n}=0&\text{ on }\Gamma_{C},\vspace{2mm}\\ \displaystyle\frac{\partial y}{\partial n}=\frac{\partial p}{\partial n}=0&\text{ on }\Gamma_{N},\\ y=p=0&\text{ on }\Gamma_{D},\\ \end{cases}$
@@
(10) $\left(\boldsymbol{x}_{1},\boldsymbol{x}_{2}\right)_{X}=\left(y_{1},y_{2}\right)_{\Omega}+\left(u_{1},u_{2}\right)_{\Gamma_{C}}\,,$
@@
(11) $\left\lVert{\boldsymbol{x}}\right\rVert_{X}^{2}=\left\lvert\left\lvert\left\lvert y\right\rvert\right\rvert\right\rvert_{Y}^{2}+\left\lVert{u}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}^{2},$
@@
(12) ${\mathcal{A}}\left(\boldsymbol{x},\boldsymbol{\varsigma}\right)=\left(y,w\right)_{\Omega_{\text{obs}}}+\alpha\left(u,v\right)_{\Gamma_{C}}\quad\forall\boldsymbol{x}=(y,u),\boldsymbol{\varsigma}=(w,v)\in X\,,$
(13) ${\mathcal{B}}\left(q,\boldsymbol{x}\right)=a(y,q)-\left(u,q\right)_{\Gamma_{C}}\quad\forall\boldsymbol{x}\in X,\,\forall q\in Y\,.$
@@
3. VEM Numerical Discretization
In this section, we introduce the VEM approximation of Problem 3.
Let $\mathcal{M}_{h}$ denote a polygonal tessellation of $\Omega$ conforming to $\Omega_{\text{obs}}$. $E$ denotes a general polygon of $\mathcal{M}_{h}$ with diameter $h_{E}$ and let $h$ be the maximum diameter over $\mathcal{M}_{h}$. Let $e$ denote an edge of $\mathcal{M}_{h}$, and $\mathcal{I}_{h}$ be the set of edges lying on $\Gamma_{C}$. We assume that $\mathcal{M}_{h}$ fulfills the following assumptions [6, 15]: $\exists\kappa>0$ such that
@@ 3. VEM Numerical Discretization
    for all $E\in\mathcal{M}_{h}$, $E$ is star-shaped with respect to a ball of radius $\rho\geq\kappa h_{E}$;
@@
    for all edges $e$ of $\mathcal{M}_{h}$, $\left\lvert{e}\right\rvert\geq\kappa h_{E}$.
@@
(16) $\forall\,w\in\mathrm{H}^{1}\!(E),\quad\left(\nabla\left(\Pi^{\nabla,E}_{k}w-w\right),\nabla z_{k}\right)_{E}=0\;\;\forall\,z_{k}\in\mathbb{P}_{k}\!\left(E\right).$
In order to define $\Pi^{\nabla,E}_{k}$ uniquely, we set $\int_{\partial E}\left(\Pi^{\nabla,E}_{k}u-u\right)\,\mathrm{d}s=0$. For any given $E\in\mathcal{M}_{h}$, let $Y_{h}(E)$ be the local Virtual Element Space:
$Y_{h}(E)=\{w_{h}\in\mathrm{H}^{1}\!(E)\colon\Delta w_{h}\in\mathbb{P}_{k}\!\left(E\right),\,w_{h_{|e}}\in\mathbb{P}_{k}\!\left(e\right)\,\forall e\subset\partial E,\,w_{h}\in C^{0}\!\left(\partial E\right),\\ \left(w_{h},z_{k}\right)_{E}=\left(\Pi^{\nabla,E}_{k}w_{h},z_{k}\right)_{E}\;\forall z_{k}\in\mathbb{P}_{k}\!\left(E\right)\setminus\mathbb{P}_{k-2}\!\left(E\right)\}\,.$
We recall that the degrees of freedom of this space are (see [1, 4])
  • The values of $v$ at the vertices of $E$,
  • If $k\geq 2$, the values at $k-1$ internal distinct points on each edge $e\subset\partial E$,
  • If $k\geq 2$, the scaled moments $\frac{1}{\left\lvert{E}\right\rvert}\int_{E}v\,m_{i}$, where $\{m_{i}\}_{i=1}^{\dim\mathbb{P}_{k-2}\!\left(E\right)}$ is a basis of $\mathbb{P}_{k-2}\!\left(E\right)$.
@@
(17) $\left\lvert\left\lvert\left\lvert w\right\rvert\right\rvert\right\rvert_{Y(E)}^{2}=\left\lVert{\sqrt{\mathcal{K}}\nabla w}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}+\left\lVert{\sqrt{\gamma}w}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2}\,,$
@@
(18) $\leavevmode\nobreak\ U_{h}(E){\;\vcentcolon=\;}\left\{u_{h}\in U:u_{h}\in\mathbb{P}_{1}\!\left(e\right)\;\forall e\in\mathcal{I}_{h}\,,u_{h}\in C^{0}\!\left(\Gamma_{C}\right)\right\},$
and it is equipped with the $\mathrm{L}^{2}$-norm on $\Gamma_{C}$. To define the discrete formulation of Problem 3, let $\Pi^{0,E}_{k-1}$ be the local $\mathrm{L}^{2}$-orthogonal projector into the space of polynomials of degree up to $k-1$, i.e., for each $w\in\mathrm{L}^{2}\!(E)$, $\int_{E}(\Pi^{0,E}_{k-1}w-w)=0$ and, for each $\boldsymbol{w}=\left(w_{1},w_{2}\right)^{\intercal}\in\left[\mathrm{L}^{2}\!\left(E\right)\right]^{2}$, $\Pi^{0,E}_{k-1}\boldsymbol{w}=\left(\Pi^{0,E}_{k-1}w_{1},\Pi^{0,E}_{k-1}w_{2}\right)^{\intercal}$. Notice that $\Pi^{\nabla,E}_{k}$ and $\Pi^{0,E}_{k-1}$ are computable for any $w_{h}\in Y_{h}(E)$ exploiting the degrees of freedom.
For each $E\in\mathcal{M}_{h}$, let ${a}^{E}_{h}:Y_{h}(E)\times Y_{h}(E)\to\mathbb{R}$ and ${m}^{E}_{h}:Y_{h}(E)\times Y_{h}(E)\to\mathbb{R}$ be the local bilinear forms defined as
(19) ${a}^{E}_{h}\left(w_{h},q_{h}\right)\vcentcolon=\left(\mathcal{K}\Pi^{0,E}_{k-1}\nabla w_{h},\Pi^{0,E}_{k-1}\nabla q_{h}\right)_{E}+\left\lVert{\mathcal{K}}\right\rVert_{\infty}S^{E}\left(w_{h}-\Pi^{\nabla,E}_{k}w_{h},q_{h}-\Pi^{\nabla,E}_{k}q_{h}\right)\\ +\left(\gamma\Pi^{0,E}_{k}w_{h},\Pi^{0,E}_{k}q_{h}\right)_{E}+\left\lVert{\gamma}\right\rVert_{\infty}\left\lvert{E}\right\rvert S^{E}\left(w_{h}-\Pi^{\nabla,E}_{k}w_{h},q_{h}-\Pi^{\nabla,E}_{k}q_{h}\right)$

### v2

# Stabilization-Free General Order Virtual Element Methods for Neumann Boundary Optimal Control Problems in Saddle Point Formulation

@@ Abstract
In this work, we explore the application of Stabilization-Free Virtual Element Methods for Neumann boundary Optimal Control Problems in saddle point formulation. The method is proposed for arbitrary polynomial order of accuracy and general polygonal meshes. Our contribution includes a rigorous a priori error estimate that holds for general polynomial order. On the numerical side, we present (i) an initial convergence test that reflects our theoretical findings, (ii) a second test analyzing the role of the stabilization term in the Virtual Element Method (VEM) formulation and its influence on the approximation error, and (iii) a third test case based on a more application-oriented experiment. The stabilization-free approach is proposed as an alternative strategy to circumvent issues related to the choice of the stabilization parameter in standard VEM formulations.
keywords:
Virtual Element Methods, Neumann boundary optimal control problems , stabilization-free , saddle point formulation
2020 MSC:
35B37 , 65N12 , 65N15 , 65N30
[…]
1 Introduction
@@
In this paper, we consider a Stabilization-Free Virtual Element Method (SFVEM) approximation for optimal Neumann boundary control. Optimal Control Problems (OCPs) governed by partial differential equations (PDEs) are widespread in many practical applications in scientific and engineering contexts, see e.g. [1, 26, 28, 29] and the references therein. The main goal of OCPs is to reach a beneficial behavior of the PDE solution by employing an external variable, the control. The role of the control consists in making the state solution similar to a desired configuration employing a functional minimization. The control is distributed when its action is defined all over the domain, while we talk about boundary control if the control affects the boundary conditions. Due to its coupled nature, the numerical analysis for OCPs governed by PDEs is a fascinating mathematical challenge. Several works have been proposed with Finite Element Methods, Discontinuous Galerkin Methods, Finite Volumes and hybrid methods (see for instance [14, 18, 37]).
@@
Lately, great attention has been paid to VEM, a family of polygonal methods for simulating PDEs. VEM can be applied to arbitrary polygonal meshes, guaranteeing huge flexibility in terms of geometrical complexity, and they might play a crucial role in many real-life scenarios. Indeed, after being conceived in the seminal paper [2], they have been applied to various contexts of interest. The investigation of boundary OCPs through VEM is still limited, yet of utmost importance. Indeed, the versatility of arbitrary polygonal meshes, combined with the action of control strategies on complex geometries, might not only be a challenging theoretical topic but also relevant from a practical viewpoint. To our knowledge, $C^{1}$-VEM is presented in [12] for distributed OCPs governed by elliptic equations and with a constrained state variable. The VEM discretization for distributed control is the topic of several investigations, see e.g. [30, 32, 33, 35]. Some VEM works related to boundary OCPs are [34] for the Dirichlet case with control constraints and [20] for the Neumann case with control constraints and approximation of the first order. A key ingredient of the standard VEM approximation for any problem of interest is the use of discrete bilinear forms composed of a consistent polynomial part and a stabilization term ensuring coercivity. It is known that the choice of stabilization remains somewhat arbitrary due to its problem dependance, and its influence on the system cannot be reliably controlled in more complex settings, such as the coupled problems and control applications considered here. In the literature, many alternatives have been proposed that do not rely on the stabilization term. Starting from the works [16, 6, 8], a research line has grown, defining the so-called Stabilization-Free VEM (SFVEM) and focusing on the definition of VEM bilinear forms that are self‑stabilized through the use of higher‑order polynomial projections (see for instance [19, 15, 7, 36, 11]). For further studies on the role of stabilization and SFVEM, the reader may refer to [27, 4]. In this contribution, we investigate the SFVEM within the framework of OCPs, employing the strategy proposed in [9, 5] which exploits the properties of divergence‑free polynomial spaces. In particular, we focus on the unconstrained Neumann boundary control problem in a saddle point structure. This naturally derives from the optimality conditions and considers the coupled nature of the system leading to more robust and stable systems, solving for all the variables at once. This formulation is generally more effective than iterative algorithms. For these kinds of coupled systems, SFVEM might represent an asset in terms of robustness of the method compared to stabilized VEM. To the best of our knowledge, we differentiate from previous contributions in the following terms:
  • the presented VEM scheme does not require any additional stabilization term and has arbitrary polynomial accuracy;
@@
  • we analyze the problem in saddle point formulation as presented in [22], hence providing a complete proof of well-posedness and a priori error estimates;
  • the numerical tests cover high-order schemes, and we numerically investigate how the stabilization parameter affects the method’s accuracy, assessing the flexibility and robustness of the proposed stabilization-free scheme.
@@
The paper is structured as follows: in Section 2 we introduce the linear-quadratic Neumann boundary OCP at the continuous level in a saddle point framework. Section 3 focuses on the Stabilization-Free Virtual Element Method (SFVEM) approximation of such optimization problem and proves the well-posedness of the saddle point structure. In Section 4 we derive optimal a priori error estimates. Section 5 presents numerical results that confirm the expected convergence behavior, we also analyze how the accuracy of the stabilized VEM formulation depends on the choice of the stabilization parameter, and we present the proposed SFVEM method as a robust strategy for a more application-oriented test.
Finally, conclusions follow in Section 6.
@@
2 Problem Formulation
Let us consider $\Omega\subset\mathbb{R}^{2}$ as a bounded domain with a polygonal boundary. We denote as $\Gamma_{D}$ and $\Gamma_{N}$ the portions of the boundary $\partial\Omega$ where Dirichlet and Neumann conditions apply. Moreover, we define a control boundary $\Gamma_{C}$, where the control acts in form of Neumann condition. We remark that $\Gamma_{D}\cup\Gamma_{N}\cup\Gamma_{C}=\partial\Omega$ and they are pairwise disjoint. Moreover, for simplicity we assume $\Gamma_{D}\neq\emptyset$.
@@ 2 Problem Formulation
The case of non-homogeneous boundary conditions can be tackled by standard lifting procedures. Given a desired state $y_{\mathrm{d}}\in Y_{\text{d}}=L^{2}(\Omega_{\text{obs}})$, with $\Omega_{\text{obs}}\subset\Omega$, a forcing term $f\in L^{2}(\Omega)$ and a penalization parameter $\alpha>0$ on the control action, we solve the following problem.
@@
[…]
with $\beta\in(L^{\infty}(\Omega))^{2}$ is divergence-free and $\gamma\in L^{\infty}(\Omega)$ is non negative on $\Omega$. Moreover, we consider the symmetric diffusive tensor $\mathcal{K}\in(L^{\infty}(\Omega))^{2\times 2}$ satisfying
$\mathcal{K}_{0}\,|\nu|^{2}\;\leq\;\nu\cdot\mathcal{K}(x)\,\nu\;\leq\;\mathcal{K}_{1}\,|\nu|^{2},\quad\forall\,\nu\in\mathbb{R}^{2},\ \forall\,x\in\Omega,$ (2)
for some positive constants $\mathcal{K}_{0}$ and $\mathcal{K}_{1}$ and $|\cdot|$ denoting the Euclidean norm.
@@
${a(y,q)}=\left(u,q\right)_{\Gamma_{C}}+\left(f,q\right)_{\Omega}\quad\forall q\in Y,$ (3)
@@
$a(y,q)\vcentcolon=\left(\mathcal{K}\nabla y,\nabla q\right)_{\Omega}+\left(\beta\nabla y,q\right)_{\Omega}+\left(\gamma y,q\right)_{\Omega}\,\quad\forall y,q\in Y.$ (4)
Since $a(\cdot,\cdot)$ is continuous and coercive and $\left(u,q\right)_{\Gamma_{C}}$ is continuous, (3) is well-posed. Let $\left\lvert\left\lvert\left\lvert\cdot\right\rvert\right\rvert\right\rvert_{Y}$ denote the following norm: for each $w\in Y$
$\left\lvert\left\lvert\left\lvert w\right\rvert\right\rvert\right\rvert_{Y}^{2}{\;\vcentcolon=\;}\left\lVert{\sqrt{\mathcal{K}}\nabla w}\right\rVert_{\mathrm{L}^{2}\!(\Omega)}^{2}.$ (5)
@@
For any $w\in Y$, there exists $\tilde{C}>0$ such that
$\left\lVert{w}\right\rVert_{\mathrm{H}^{1}\!(\Omega)}\coloneqq\left\lVert{\nabla w}\right\rVert_{\mathrm{L}^{2}\!(\Omega)}^{2}+\left\lVert{w}\right\rVert_{\mathrm{L}^{2}\!(\Omega)}^{2}\leq\tilde{C}\left\lvert\left\lvert\left\lvert w\right\rvert\right\rvert\right\rvert_{Y}$ (6)
holds true, due to a Poincaré inequality and the regularity hypothesis of $\mathcal{K}$. Thanks to the same hypothesis, we have that, for a positive constant $C$
$\left\lVert{w}\right\rVert_{\mathrm{L}^{2}\!(\Omega)}^{2}\leq C\left\lvert\left\lvert\left\lvert w\right\rvert\right\rvert\right\rvert_{Y}.$ (7)
holds true.
@@
$\mathscr{L}(y,u,p)=J(y,u)+\displaystyle a(y,p)-\left(u,p\right)_{\Gamma_{C}}-\left(f,p\right)_{\Omega}.$ (9)
@@
[…]
where $D_{y},D_{u}$ and $D_{p}$ denotes the differentiation with respect to the three variables, respectively. In this way, the PDE-constrained optimization problem is reframed in an unconstrained optimization and the minimum of (8) is represented by the state and control components of the stationary point of (9). The strong formulation of the optimality system reads:
@@
[…]
@@
$\left(\boldsymbol{x}_{1},\boldsymbol{x}_{2}\right)_{X}=\left(y_{1},y_{2}\right)_{\Omega}+\left(u_{1},u_{2}\right)_{\Gamma_{C}}\,$, (12)
@@
$\left\lVert{\boldsymbol{x}}\right\rVert_{X}^{2}=\left\lvert\left\lvert\left\lvert y\right\rvert\right\rvert\right\rvert_{Y}^{2}+\left\lVert{u}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}^{2},$ (13)
@@
${\mathcal{A}}\left(\boldsymbol{x},\boldsymbol{\varsigma}\right)=\left(y,w\right)_{\Omega_{\text{obs}}}+\alpha\left(u,v\right)_{\Gamma_{C}}\quad\forall\boldsymbol{x}=(y,u),\boldsymbol{\varsigma}=(w,v)\in X\,$, (14)
${\mathcal{B}}\left(q,\boldsymbol{x}\right)=a(y,q)-\left(u,q\right)_{\Gamma_{C}}\quad\forall\boldsymbol{x}\in X,\,\forall q\in Y\,$, (15)
@@
3 Numerical Discretization
In this section, we introduce the Stabilization-Free Virtual Element (SFVEM) approximation of Problem 3 following [9, 5]. Let $\mathcal{M}_{h}$ denote a polygonal tessellation of $\Omega$ conforming to $\Omega_{\text{obs}}$. $E$ denotes a general polygon of $\mathcal{M}_{h}$ with diameter $h_{E}$ and let $h$ be the maximum diameter over $\mathcal{M}_{h}$. Let $e$ denote an edge of $\mathcal{M}_{h}$, $\mathcal{I}_{h}$ be the set of edges lying on $\Gamma_{C}$ and $\mathcal{M}_{h,\text{obs}}$ be the set of polygons $E$ in $\Omega_{\text{obs}}$. We assume that $\mathcal{M}_{h}$ fulfills the following assumptions: $\exists\kappa>0$ such that
@@ 3 Numerical Discretization
    for all $\in\mathcal{M}_{h}, EE$ is star-shaped with respect to a ball of radius $\rho\geq\kappa h_{E}$;
@@
    for all $E\in\mathcal{M}_{h}$, for all edges $e$ of $\mathcal{M}_{h}$, $\left\lvert{e}\right\rvert\geq\kappa h_{E}$.
@@
$\forall\,w\in\mathrm{H}^{1}\!(E),\quad\left(\nabla\left(\Pi^{\nabla,E}_{k}w-w\right),\nabla\pi_{k}\right)_{E}=0\;\;\forall\,\pi_{k}\in\mathbb{P}_{k}\!\left(E\right).$ (18)
In order to define $\Pi^{\nabla,E}_{k}$ uniquely, we set $\int_{\partial E}\left(\Pi^{\nabla,E}_{k}w-w\right)\,\mathrm{d}s=0$ if $k=1$ and $\int_{E}\left(\Pi^{\nabla,E}_{k}w-w\right)\,\mathrm{d}s=0$ otherwise. For any given $E\in\mathcal{M}_{h}$, let $Y_{h}(E)$ be the local Virtual Element Space:
$Y_{h}(E)=\{w_{h}\in\mathrm{H}^{1}\!(E)\colon\Delta w_{h}\in\mathbb{P}_{k}\!\left(E\right),\,w_{h_{|e}}\in\mathbb{P}_{k}\!\left(e\right)\,\forall e\subset\partial E,\,w_{h}\in C^{0}\!\left(\partial E\right),\\ \left(w_{h},\pi_{k}\right)_{E}=\left(\Pi^{\nabla,E}_{k}w_{h},\pi_{k}\right)_{E}\;\forall\pi_{k}\in\mathbb{P}_{k}\!\left(E\right)/\mathbb{P}_{k-2}\!\left(E\right)\}\,.$
We recall that the degrees of freedom of this space are (see [3])
  • The values of $w_{h}$ at the vertices of $E$,
  • If $k\geq 2$, the values at k1k-1 internal distinct points on each edge $e\subset\partial E$,
  • If $k\geq 2$, the scaled moments $\frac{1}{\left\lvert{E}\right\rvert}\int_{E}w_{h}\,m_{i}$, where $\{m_{i}\}_{i=1}^{\dim\mathbb{P}_{k-2}\!\left(E\right)}$ is a basis of $\mathbb{P}_{k-2}\!\left(E\right)$.
@@
$\left\lvert\left\lvert\left\lvert w\right\rvert\right\rvert\right\rvert_{Y(E)}^{2}=\left\lVert{\sqrt{\mathcal{K}}\nabla w}\right\rVert_{\mathrm{L}^{2}\!(E)}^{2},$ (19)
@@
$\penalty 10000\ U_{h}(E){\;\vcentcolon=\;}\left\{v_{h}\in U:v_{h}\in\mathbb{P}_{1}\!\left(e\right)\;\forall e\in\mathcal{I}_{h}\,,v_{h}\in C^{0}\!\left(\Gamma_{C}\right)\right\},$ (20)
and it is equipped with the $\mathrm{L}^{2}$-norm on $\Gamma_{C}$. We now focus on the ingredients of the SFVEM approximation of the bilinear forms of Problem 3.
Let $\Pi^{0,E}_{k-1}$ be the local $\mathrm{L}^{2}$-orthogonal projector into the space of polynomials of degree up to k1k-1, i.e. for each $w\in\mathrm{L}^{2}\!(E)$, $\int_{E}(\Pi^{0,E}_{k-1}w-w)\pi_{k-1}=0$ for each $\pi_{k-1}\in\mathbb{P}_{k-1}\!\left(E\right)$. Then, for each $\boldsymbol{w}=\left(w_{1},w_{2}\right)^{\intercal}\in\left[\mathrm{L}^{2}\!\left(E\right)\right]^{2}$, we define the vector-valued projection operator component-wise, e.g. $\Pi^{0,E}_{k-1}\boldsymbol{w}=\left(\Pi^{0,E}_{k-1}w_{1},\Pi^{0,E}_{k-1}w_{2}\right)^{\intercal}$.
Moreover, under assumptions A.1 and A.2, for each $E\in\mathcal{M}_{h}$ and for $\ell_{E}\in\mathbb{N}$, we define the space
$\mathcal{P}_{k,\ell_{E}}\vcentcolon=\bigl[P_{k-1}(E)\bigr]^{2}\;\oplus\;\boldsymbol{\mathrm{curl}}\,\Bigl(P_{k+\ell_{E}}(E)\setminus P_{k}(E)\Bigr)=\mathbf{x}\,P_{k-2}(E)\;\oplus\;\boldsymbol{\mathrm{curl}}\,\,P_{k+\ell_{E}}(E),$ (21)
with $\boldsymbol{\mathrm{curl}}\,\pi=\left(\frac{\partial\pi}{\partial y},-\frac{\partial\pi}{\partial x}\right)$ for any $\pi\in\mathbb{P}_{\ell_{E}+1}\!\left(E\right)$. Let us define $\Pi^{0,E}_{\mathcal{P}}\nabla:H^{1}(E)\rightarrow\mathcal{P}_{k,\ell_{E}}$ the projection operator defined by the following orthogonality condition for any given $w$
$\left(\Pi^{0,E}_{\mathcal{P}}\nabla w,\pi\right)_{E}=\left(\nabla w,\pi\right)_{E}\quad\forall\pi\in\mathcal{P}_{k,\ell_{E}}.$ (22)
Notice that for $w_{h}\in Y_{h}(E)$, the polynomial projections $\Pi^{\nabla,E}_{k}w_{h}$, $\Pi^{0,E}_{k}w_{h}$, $\Pi^{0,E}_{k-1}\nabla w_{h}$ and $\Pi^{0,E}_{\mathcal{P}}\nabla w_{h}$ are computable exploiting only the degrees of freedom of $w_{h}$.
Let ${a}^{E}_{h}:\mathrm{H}^{1}\!(E)\times\mathrm{H}^{1}\!(E)\to\mathbb{R}$ and ${m}^{E}_{h}:\mathrm{H}^{1}\!(E)\times\mathrm{H}^{1}\!(E)\to\mathbb{R}$ be the local bilinear forms defined as, for all $w_{h},q_{h}\in\mathrm{H}^{1}\!(E)$
${a}^{E}_{h}\left(w_{h},q_{h}\right)$ $\vcentcolon=\left(\mathcal{K}\Pi^{0,E}_{\mathcal{P}}\nabla w_{h},\Pi^{0,E}_{\mathcal{P}}\nabla q_{h}\right)_{E}$ (23)
$\qquad\qquad+\left(\beta\Pi^{0,E}_{k-1}\nabla w_{h},\Pi^{0,E}_{k}q_{h}\right)_{E}+\left(\gamma\Pi^{0,E}_{k}w_{h},\Pi^{0,E}_{k}q_{h}\right)_{E},$